# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ Appendix A Proofs and Auxiliary Results
We now employ the expressions deduced in (66) in order to prove the recursive feasibility of the problems formulated in (42). To do so, choose any desired initial time instant $k_{0}$ for the optimization problem associated with the $i^{\text{th}}$ area index. For all $t\in\{1:\rho_{i}\}$, consider any $\widetilde{\theta}_{xi}[k_{0}+t]\in\widetilde{\Theta}_{xit}$ and any $\widetilde{\theta}_{ui}[k_{0}+t]\in\widetilde{\Theta}_{uit}$ which may arise in the context of the aforementioned optimization problem, in order to define $\widetilde{\theta}_{xit}:=\widetilde{\theta}_{xi}[k_{0}+t]$ along with $\widetilde{\theta}_{u\ell t}:=\mathrm{row}_{(\ell-\alpha_{ui})}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)$, with $\ell\in\mathcal{A}_{ui}$. Due to the arguments made in the previous paragraph of this proof, it follows that there must exist some appropriately constrained vectors denoted $u_{1it}$ and $u_{2it}$ along with $p_{xit}$ and $p_{u\ell t}$, with $t\in\{1:\rho_{i}\}$ and $\ell\in\mathcal{A}_{ui}$, such that the identities from (66) hold. Therefore, by setting $T_{i}=\rho_{i}$ and by selecting an arbitrary $\overline{T}_{i}\in\mathbb{N}$, it is always possible to choose the command signal vectors which appear in (42) as $u_{s1i}[k_{0}+t-1]:=u_{1ij}\in\mathcal{U}_{s1i}$ and as $u_{s2i}[k_{0}+t-1]:=u_{2ij}\in\mathcal{U}_{s2i}$, for all $t\in\{1:T_{i}\}$. A direct consequence of all these choices is the fact that
$\xi_{i}[k_{0}+t]=\sum_{j=1}^{t}A_{si}^{(j-1)}(B_{s1i}u_{1ij}+B_{s2i}u_{2ij}),\vspace{-3mm}$ (67)
@@
At this point in the proof, we have shown that by setting $T_{i}=\rho_{i}$ and by choosing an arbitrary $\overline{T}_{i}\in\mathbb{N}$, there exists at least one sequence of command vectors which satisfies the constraints of the problems formulated in (42), thus rendering them feasible. Recall, moreover, that the time instant $k_{0}$ along with the area index $i\in\{1:N\}$ were chosen at random at the beginning of the proof, which makes the feasibility of all the aforementioned problems a generic property. By recalling now Assumption A1), we employ Proposition 6.1 to compute the second-layer command values at time instant $k=k_{0}$ and, by iterating the procedure detailed in Algorithm 1, we arrive at time instant $k=k_{0}+1$, in which Assumption A1) is still valid and in which all $N$ local optimization problems remain feasible. It follows that this procedure may be successfully repeated at any subsequent instant $k\geq k_{0}$, thus guaranteeing recursive feasibility, i.e., point $i)$ of our result, while also ensuring the satisfaction of the constraints which make up point $ii)$ of our result.

### v4

@@ Appendix B Proofs and Auxiliary Results
for all $t\in\{1:\rho_{i}\}$ and all $\ell\in\mathcal{A}_{ui}$, where $\widetilde{\theta}_{xit}\in\widetilde{\Theta}_{xit}$, $\widetilde{\theta}_{u\ell t}\in\left(e_{(\ell-\alpha_{ui})}^{\top}\widetilde{\Theta}_{uit}\right)$, ${p}_{xit}\in\mathcal{P}_{1it}$ and ${p}_{u\ell t}\in\mathcal{P}_{2\ell t}$. We now employ the expressions deduced in (64) in order to prove the recursive feasibility of the problems formulated in (8). To do so, choose any desired initial time instant $k_{0}$ for the optimisation problem associated with the $i^{\text{th}}$ area index. For all $t\in\{1:\rho_{i}\}$, consider any $\widetilde{\theta}_{xi}[k_{0}+t]\in\widetilde{\Theta}_{xit}$ and any $\widetilde{\theta}_{ui}[k_{0}+t]\in\widetilde{\Theta}_{uit}$ which may arise in the context of the aforementioned optimisation problem, in order to define $\widetilde{\theta}_{xit}:=\widetilde{\theta}_{xi}[k_{0}+t]$ along with $\widetilde{\theta}_{u\ell t}:=\mathrm{row}_{(\ell-\alpha_{ui})}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)$, with $\ell\in\mathcal{A}_{ui}$. Due to the arguments made in the previous paragraph of this proof, it follows that there must exist some appropriately constrained vectors denoted $u_{1it}$ and $u_{2it}$ along with $p_{xit}$ and $p_{u\ell t}$, with $t\in\{1:\rho_{i}\}$ and $\ell\in\mathcal{A}_{ui}$, such that the identities from (64) hold. Therefore, by setting $T_{i}=\rho_{i}$ and by selecting an arbitrary $\overline{T}_{i}\in\mathbb{N}$, it is always possible to choose the command signal vectors which appear in (8) as $u_{s1i}[k_{0}+t-1]:=u_{1ij}\in\mathcal{U}_{s1i}$ and as $u_{s2i}[k_{0}+t-1]:=u_{2ij}\in\mathcal{U}_{s2i}$, for all $t\in\{1:T_{i}\}$. A direct consequence of all these choices is the fact that
$\xi_{i}[k_{0}+t]=\sum_{j=1}^{t}A_{si}^{(j-1)}(B_{s1i}u_{1ij}+B_{s2i}u_{2ij}),\vskip-8.53581pt$ (65)
@@
At this point in the proof, we have shown that by setting $T_{i}=\rho_{i}$ and by choosing an arbitrary $\overline{T}_{i}\in\mathbb{N}$, there exists at least one sequence of command vectors which satisfies the constraints of the problems formulated in (8), thus rendering them feasible. Moreover, recall that the time instant $k_{0}$ and the area index $i\in\{1:N\}$ were chosen arbitrarily at the beginning of the proof, thus making the feasibility of all the aforementioned problems a generic property. By recalling now Assumption A1), we employ Proposition 6.1 to compute the second-layer command values at time instant $k=k_{0}$ and, by iterating the procedure detailed in Algorithm 1, we arrive at time instant $k=k_{0}+1$, in which Assumption A1) is still valid and in which all $N$ local optimisation problems remain feasible. It follows that this procedure may be successfully repeated at any subsequent instant $k\geq k_{0}$, thus guaranteeing recursive feasibility, i.e., point i) of our result, while also ensuring the satisfaction of the constraints which make up point ii) of our result.